OpenStudy (anonymous):

log(subscript2)(x-3)=5

OpenStudy (anonymous):

\[\huge \log_{2}(x-3)=5\]

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Now you can take the two in the subscript to the other side like this \[\huge (x-3)=2^{5}\]

OpenStudy (anonymous):

dont you have to distribute?

OpenStudy (anonymous):

the log

OpenStudy (anonymous):

See log is nothing but in words it reads as what power 2 must be raised to get x-3 = 5 so 2 must be raised to the 5 th power to get x-3 and that is what i have written

OpenStudy (anonymous):

alright i get it thank you

OpenStudy (anonymous):

i ahd that before lol

OpenStudy (anonymous):

please continue dude

OpenStudy (anonymous):

what will u do after that

OpenStudy (anonymous):

oh x=35

OpenStudy (anonymous):

genius bro

OpenStudy (anonymous):

lol ikr

OpenStudy (anonymous):

hehe